OpenStudy (anonymous):

solve the equation: 3v^2=-v I got the v=0 but I do not understand how 3v^2 when solved for = (-1/3)

jimthompson5910 (jim_thompson5910):

3v^2=-v 3v^2+v = 0 v(3v+1) = 0 v = 0 or 3v+1 = 0 v = 0 or v = -1/3

OpenStudy (anonymous):

thanks